archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    net Attached to Project Dedicated servers Task Type Incident Category RBX3 Status CLOSED Percent Complete Details We have detected an isuue on this server We are working on it Date Tuesday 13 October 2015 17 18PM Reason for closing Done

    Original URL path: http://status.ovh.es/?do=details&id=11061 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Orange and OVH 3 on copper and 1 on on the optical fiber The optical fiber interconnection is UP and those of copper were vandalized There is a lack of capacity 30 to take all incoming calls By 5 30 pm the number of calls will be low enough to be cashed in degraded mode Orange is operating on site in order to to replace the copper pairs necessary for interconnection It will take a few hours In parallel a dual IP interconnection between Orange and OVH is being created one interconnection towards TH2 and the other towards GSW Comment by OVH Tuesday 13 October 2015 16 45PM The call peak falls the number of calls is low enough to be cashed The operator is still working on site Comment by OVH Tuesday 13 October 2015 16 46PM Orange has yet to work tomorrow morning to return the situation back to normal The current capacity is more than enough to keep the load of the weekend Comment by OVH Tuesday 13 October 2015 16 47PM We are expecting the operator s feedback on the work progress Comment by OVH Tuesday 13 October 2015 17 00PM We have 4 links to the 3 different Orange infrastructures 3 PRO in Paris every POP is independent and provides routing calls to our infrastructure When a link is DOWN the Orange PRO should return the call to another PRO and so on until the call ends on the PRO that is connected to our infrastructure We have noticed that this is not working well and some calls are not successful It is possible that the call arrives on one of the three PRO when it is DOWN so it returns the call to another PRO If this 2nd PRO is also DOWN and then

    Original URL path: http://status.ovh.es/?do=details&id=11058 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    11056 FS 14971 filerz2005 Attached to Project Hosting Task Type Incident Category 300gp Status CLOSED Percent Complete Details We have detected a load issue Reboot is in progress Date Tuesday 13 October 2015 16 15PM Reason for closing Done Comments

    Original URL path: http://status.ovh.es/?do=details&id=11056 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    storage incident Migration to a new datastore is in progress The VPS could be stopped for migration vps11007 vps11702 vps11651 vps11968 vps13480 vps19120 vps17575 vps11779 vps18858 vps16387 vps13912 vps11943 vps10632 vps18653 vps13196 vps10838 vps10678 vps16589 vps11215 vps13486 vps11244 vps18752 vps10982 vps10966 vps13454 vps13858 vps13491 vps14238 vps11969 vps18370 vps14417 vps11877 vps11184 vps10288 vps10279 vps11710 vps18057 vps13198 vps13648 vps14440 vps13872 vps16577 vps13873 vps17425 vps13809 vps11644 vps13588 vps10248 vps13608 vps13850 vps11831 vps11905 vps11172 vps14296 vps11527 vps17617 vps14206 vps13393 vps18384 vps13875 vps11056 vps10649 vps11648 vps10617 vps10644 vps18609 vps11556 vps13792 vps11001 vps10721 vps14427 vps11463 vps13612 vps13493 vps17807 vps13720 vps10856 vps10886 vps11650 vps17839 vps10486 vps11617 vps11289 vps16649 vps11688 vps11205 vps11768 vps18724 vps14195 vps11947 vps10215 vps17946 vps14006 vps10906 vps17849 vps16562 vps10446 vps13529 vps11981 vps16533 vps11783 vps11743 vps11707 vps18094 vps11196 vps13719 vps13880 vps16390 vps14661 vps11227 vps11503 vps18765 vps11108 vps13689 vps14063 vps10467 vps13672 vps11894 vps14375 vps14062 vps14136 vps14665 vps17660 vps13963 vps11728 vps14348 vps16348 vps11810 vps11418 vps14677 vps13942 vps16619 vps14032 vps13835 vps11612 vps17110 vps13424 vps11874 vps18080 vps11482 vps11360 vps14341 vps18648 vps14107 vps10801 vps10846 vps17856 vps13218 vps14093 vps10677 vps17851 vps13638 vps18766 vps14655 Comment by OVH Tuesday 13 October 2015 16 07PM The following VPS have been migrated to a new datastore vps11947 vps13486 vps13809 vps16619 vps17849 vps18765 vps10215 vps11289 vps13880 vps17617

    Original URL path: http://status.ovh.es/?do=details&id=11055 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Attached to Project Hosting Task Type Incident Category 60gp Status CLOSED Percent Complete Details Access loss to filerz Reboot in progress Date Tuesday 13 October 2015 15 50PM Reason for

    Original URL path: http://status.ovh.es/?do=details&id=11054 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Cloud Task Type Incident Category Rbx2a Status In progress Percent Complete Details We have detected an IO issue Verification is in progress Comments 1 Comment by OVH Tuesday 13 October 2015 15 48PM The service is operacional There is a

    Original URL path: http://status.ovh.es/?do=details&id=11053 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Attached to Project Dedicated servers Task Type Incident Category RBX2 Status CLOSED Percent Complete Details We have detected a problem on this server We are working on it Date Tuesday 13 October 2015 15 44PM Reason for closing Done Additional

    Original URL path: http://status.ovh.es/?do=details&id=11052 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    net Attached to Project Hosting Task Type Incident Category 20gp Status CLOSED Percent Complete Details We have detected a problem on this server We are working on it Date Tuesday 13 October 2015 15 41PM Reason for closing Done Additional

    Original URL path: http://status.ovh.es/?do=details&id=11051 (2015-10-15)
    Open archived version from archive



  •